OpenStudy (anonymous):

How do I work the problem out! What is 5% of 80?

OpenStudy (anonymous):

let's try this: 5% is a basically a fraction, 5/100. You could do the multiplication of \[5% * 80\] as 5/100*80 => (5*80)/100 = 400/100 which can simplify to 4/1. Or you can think that it takes 20 5's to add up to 100. so 5% is also 1/20th, so what is 1/20th of 80? What value do we need so that if we have 20 of this number (x) it will add up to 80? 20 4's add up to 80, so just one of these, 1/20th, is 4.

OpenStudy (anonymous):

hope this helped...good luck with your other problems!

OpenStudy (anonymous):

5%of 80 = .05/80=

OpenStudy (anonymous):

not 5/100 think smaller 5 /10 then simplify=1/2 so .05x 80